What is the probability that she will roll a number other than 1?
On a single toss of a six-sided die. The chance of rolling a one is 1/6, and the chance of not rolling a one is (1 - 1/6 = 5/6), or 83.333 %
Neither. The marble has higher density than breadcrumbs or croutons, so it will sink, not rise.
There are 26 letters in the English alphabet, of which 5 are vowels (a, e, i, o, u). Since each card is identical and there are 26 cards in the box, the probability of selecting any one card is 1/26.
To find the probability of selecting two cards with vowels printed, we need to use the concept of conditional probability. The probability of selecting a vowel on the first draw is 5/26, since there are 5 vowels in the box. After the first card is selected and not replaced, there are only 25 cards left in the box, of which 4 are vowels (since one vowel has already been selected). Therefore, the probability of selecting a vowel on the second draw, given that a vowel was selected on the first draw, is 4/25.
To find the probability of both events occurring (i.e., selecting two cards with vowels printed), we need to multiply the probabilities of each event together, since they are independent:
P(selecting a vowel on the first draw) x P(selecting a vowel on the second draw, given that a vowel was selected on the first draw) = (5/26) x (4/25)
Simplifying this expression, we get:
(5/26) x (4/25) = 20/650 = 2/65
Therefore, the probability of selecting two cards with vowels printed is 2/65
Gives a standardized unit that tells how far away each measurement is from the mean?
gives a standardized unit that tells how far away each measurement is from the mean
How many Indian heads are there?
3, Indian Head cent, Indian Head nickel, Indian Head Gold half eagle
The probability is approx 0.3528
When rolling a pair of six sided cubes what is the chance of rolling an eight?
Assuming fair dice,
each of the pair has an equal chance of landing on any of their six faces.
These events should occur independently, so there are 36 total possibilities of equal probability.
That is, each possibility has a probability 1/36.
An 8 can be made with the possibilities (2,6), (3,5), (4,4), (5,3), (6,2).
Any of these possibilities will work so we want the probability that any of them will occur.
This can be calculated by adding their probabilities since they are mutually exclusive events.
There are 5 of them, each with probability 1/36. So the probability of rolling an 8 is therefore
5/36
What is the chance of not getting an ace from a pack of cards?
The chances are 12 out of 13 that you will not get an ace on the first draw.
What are the odds of throwing a six on a dice two times?
If you roll the die often enough, it is certainty.
On just two rolls and if the die is fair, the probability is 1/36.
Do identical twins have to be the same sex?
Yes. Identical twins come from one fertilized egg that has been split in two. Therefore, whatever sex the single egg would've been, the twins will be.
You can draw any one of 80 different marbles.
Any one of 50 draws will be successful.
Probability of success = 50/80 = 0.625 = 62.5 %
Who are two celebrities that share the same birthday?
I bet there are millions more out there who share the same, but for two that I know, it's Chris Rock and Ashton Kutcher on February 7.
What will be the probable outcome of the town's new policy?
The town will save about 44 million gallons of water each year.
You've described a couple of events, but we don't know the question.
That's like asking: Two boys and a girl got on a bus. What is the answer ?
He can and it is not.
In a standard deck of 52 playing cards what is the probability of drawing two ace on one draw?
0.45% (4 in 52) multiplied by (3 in 51) then multiplied by 100 to make a percentage ;)
When Two coins are tossed what is the probability of a or b happening?
This question can be rather easily answered,
as soon as outcomes 'a' and 'b' are defined.
Pr(a sum ≤ 5 at least once in 3 rolls of 2 dice)
= 1 - Pr(a sum ≤ 5 never appears in 3 rolls of 2 dice)
= 1 - Pr(a sum ≤ 5 does not appear in a roll of 2 dice)3
= 1 - Pr(a sum > 5 appears in a roll of 2 dice)3 = 1 - (26/36)3
= 1 - (13/18)3 = 1 - 2197/5832 = 3635/5832 = 62.3% approx.